OpenStudy (anonymous):

The area of a rectangular field is given by the triomial t²-4t-45. THe length of the rectangle is t+5. What is the expression for the width of the field?

OpenStudy (whpalmer4):

The area of a rectangle is A = l*w If we do a bit of trivial algebra, we can rearrange that to be \[w = \frac{A}{l}\] We know \(A = t^2-4t-45\) and \(l = t+5\) So if we plug in our values for \(A\) and \(l\) we get \[w = \frac{t^2-4t-45}{t+5}\] At this point, either factor the numerator and hope something cancels out, or do the polynomial division by long division or synthetic division.

Directrix (directrix):

@JaneDoe100 Factor t^2 - 4t - 45 and post what you get here. Okay?

OpenStudy (anonymous):

I dont think that can be factored since there are no like terms

Directrix (directrix):

What are two numbers that multiply to a*c = 1*(-45) = -45 AND add to b= -4 ? The expression t^2 - 4t - 45 will factor.

Directrix (directrix):

@JaneDoe100 What about those numbers?

OpenStudy (anonymous):

-45 -5 -4 -1 1 4 5 45

OpenStudy (anonymous):

these are the factors but in there pairs none of them add to -4

Directrix (directrix):

How about -9 and +5 ?

Directrix (directrix):

t^2 - 4t - 45 = t^2 - 9t + 5t -45 = t(t - 9) + 5 (t - 9) = (t - 9) (t + 5)

Directrix (directrix):

Area of Rectangle = Length times Width t²-4t-45 = (t + 5) W (t - 9) (t + 5) = (t + 5) W [(t - 9) (t + 5)] / (t +5) = W (t - 9) = W --> @JaneDoe100

OpenStudy (whpalmer4):

To do this by straight division: t -9 ------------------- t+5 | t^2 - 4t -45 - ( t^2 +5t ) <- "t" is biggest value we can multiply with (t+5) --------- -9t - 45 - (-9t - 45) <- 9 is biggest value we can multiply with (t+5) ------------ 0 W = t-9